# QMetro++ - Python optimization package for large scale quantum metrology with customized strategy structures

### v1

@@ 45 Tsr = x * ket_bra(
49# If not provided derivative is by
50# default 0:
51T = ParamChannel(
@@
53)
54
55Lambda_tilde = Phi.link_env(T)
56# or equivalently:
57Lambda_tilde = Phi * T
@@
4
5N = 3
6d_a = 2
7# QFI for adaptive startegy and
8# correlated noise
9qfi, *_ = iss_tnet_adaptive_qfi(Lambda_tilde, N, d_a)
10
11# Input state of the environment is by
12# default a maximally mixed state but
13# can be specified using env_inp_state
14# parameter:
15rhoE = identity(2) / 2
16qfi, *_ = iss_tnet_adaptive_qfi(Lambda_tilde, N, d_a, env_inp_state=rhoE)
Bounds in this case can be computed using ad_bounds_correlated and ad_asym_bounds_correlated. In contrast to the bounds for uncorrelated case these bounds are not necessarily tight asymptotically. They are computed by upper-bounding QFI for adaptive strategy where every $m$-th control operation has access to the environment, for some integer $m$. This leads to a leakage of information form the environment, which may make the bound less tight. Naturally, the bigger $m$ makes the bound tighter but at the same time it exponentially increases the execution time. Realistically, one can compute the bound for $m\lesssim 5$ in case of qubit channels. An exemplary code to compute the bound is given below:
@@
3N = 10
4m = 3
5ns, bound = ad_bounds_correlated(Lambda_tilde, N, m)
where ns $=[1,1+m,1+2m,...,]$, bound$=[b_{1},b_{1+m},b_{1+2m},...]$ and $b_{n}$ is an upper bound on $F_{Q}^{(n)}(\Lambda_{\theta})$.
Note, that the construction (51), (52) is not unique and correlations (50) can be modeled using different $\tilde{\Lambda}_{\theta}$. For example a quantum channel:
$\mathrm{t}:\mathcal{L}(\mathcal{E})\ni\rho\mapsto\sum_{s,r}t_{sr}\rho\,t_{sr}^{\dagger}\in\mathcal{L}(\mathcal{E}),\\ t_{sr}=\sqrt{\frac{1+sr\sqrt{c}}{2}}\ket{s}\bra{r},$ (53)
@@
Values of the QFI for anti-correlated ($c=-0.75$) dephasing case are shown in Figs.7 (B) and (C). Note, that the presence of (anti)correlations significantly increases the obtained QFI which becomes even larger then the upper-bounds for uncorrelated case. Interestingly, we can see that the adaptive strategy with $d_{\mathcal{A}}=2$ is optimal for $N\leq 5$, unlike in the uncorrelated dephasing case, see Fig. 6(B).
@@ 4 Advanced package usage—optimization of strategies with arbitrary structures
Figure 9: Diagram of strategy with arbitrary structure that can be optimized using iss_opt. Input state - $\rho_{0}$ might consist of MPSs and/or states on a full Hilbert. The input states goes through an arbitrary arrangement of parametrized channels, control operations and combs. The only requirements are that operations outputs are connected to operations inputs and that there are no cycles. Finally, the state obtained after application of each operation - $\rho_{\theta}$ is being measured. The measurement is specified by the form of the pre-SLD matrix $\mathfrak{L}$ which can be a MPO and/or a matrix on a full Hilbert space. States, control operations and combs can be marked as variable or constant, pre-SLD must be variable and parametrized channels must be constant.
In addition to the functions presented in Sec. 3, QMetro++ allows the user to define their own strategy and then optimize it using the ISS algorithm. Strategies are defined via a straightforward-to-use symbolic programming. User creates a tensor network representing their strategy where some of the nodes are marked as constants and others are marked as variables to be optimized over—let us call their set $\mathcal{V}$. This strategy is then plugged into function iss_opt which based on the provided data constructs pre-QFI:
@@
In Sec. 4.1 we explain the class of QMetro++ tensors (placed in qmtensor folder, see Fig. 2) which are the building blocks of strategies. Then in Sec. 4.2 we show how such strategy can be created using the example of collisional metrological model.
@@ 4.1 QMetro++ tensors
  • parameter encoding channels
@@
connected by their physical and bond spaces like in diagram in Fig. 9. Some input states, control operations, quantum combs and all measurements are marked as variables of the pre-QFI function (54) and they will be optimized by the ISS algorithm. Each non-measurement node is represented using their tensor representation (see detailed explanation in the Appendix B) and measurements are represented via a tensor representation of the pre-SLD matrix $\mathfrak{L}$.
Figure 10: UML diagram of classes used for creation of strategies with their most important attributes and methods. Subclasses inherit all attributes and methods of their parent class. Instances of GeneralTensor and its subclasses can access a SpaceDict instance through their sdict attribute. Two tensors can be contracted with each other only when they have the same space dictionaries as sdict.
In practice, this means that each node of the network has to be initialized as an object of QMetro++ tensor class (see Fig. 10). There are five kinds of tensor classes:
@@
  • ParamTensor - a subclass of ConstTensor which has an additional attribute defining its derivative over the measured parameter $\theta$. It is meant to represent parameter encoding channels $\Lambda_{\theta}$.
  • VarTensor - class of tensors which are being optimized by the ISS algorithm. Members of this class store information specifying how they should be initialized and optimized (for example that they are CPTP channels, quantum combs or pre-SLD matrices) but they do not have any value, that is they are only symbols.
@@
10# Calling sd[’a’] will return dimension
11# of space ’a’.
@@
15sd.set_bond(’x’, 2)
16print(sd[’x’]) # Prints 2.
@@
20print(sd.irange[’x’]) # Prints 2.
Note that because physical index of a tensor representations of a CJ matrix is created by joining two indices related to the appropriate physical space its range is equal to the dimension of this space squared (see details in Appendix B).
@@
Variable tensors require only spaces and their type is derived from additional attributes:
@@
14# > element of density matrix of MPS:
15vt1 = VarTensor(
@@
18)
19# > CPTP map of a channel ’a’->’b’:
20vt1 = VarTensor(
@@
3ch = par_dephasing(0.75)
4choi = ch.choi()
@@
While contraction of two ConstTensors simply creates a new ConstTensor this is not the case when one of the contracted tensor is a VarTensor because it does not have any value. Contraction with a VarTensor creates a TensorNetwork consisting of the two tensors:
@@
Simmilarly, a contraction with a TensorNetwork creates another TensorNetwork. Results of all possible contraction combinations are presented in Table 2.
@@
Table 2: Type of a contraction result for various tensor pairs: GT–GeneralTensor, CT–ConstTensor, PT–ParamTensor, VT– VarTensor, TN–TensorNetwork. Contraction with generalized tensor raises an error because instances of this class are not meant to be used in calculations. When ParamTensor is contracted with a ConstTensor or a ParamTensor the derivative of the result is computed using the Leibnitz (chain) rule.
@@ 4.2 Collisional strategy
Figure 11: (A) Diagram of collisional strategy with multiple parametrized channels $\Lambda_{\theta}$ and ancillas $\mathcal{A}_{i}$. Values of QFI normalized by number of channel uses - $N$ for: (B) dephasing ($p=0.75$) (C) amplitude damping ($p=0.75$) and different strategies: adaptive - with $d_{\mathcal{A}}=2$ blue dashed line, parallel with $d_{\mathcal{A}}=2,\;r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}=2$ red dotted line, collisional with $d_{\mathcal{A}}=2,\;r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}=2$ green dash-dotted line.
We will describe the creation of a custom strategy using the example of a collisional strategy depicted in Fig. 11(A). Physically, we can imagine this scheme to represent a situation where multiple-particles, potentially entangled with each other initially, approach one by one the sensing system with which they interact locally via control operations $C_{i}$, carrying away information about the parameter while at the same time causing some back-action on the sensing system itself.
This strategy may be viewed as a hybrid of parallel and adaptive strategy where each control operation $\mathrm{C}_{i}$ has their own ancilla which is measured immediately after the action of $\mathrm{C}_{i}$. Notice, that if all $\mathrm{C}_{i}$ were put to SWAP gates one would regain the parallel strategy. On the other hand, this strategy does not allow control operations to communicate with each other like in the most general adaptive strategy. Therefore the QFI for this strategy must be at least as good as for the parallel strategy but in general will be worse than for the adaptive strategy with ancilla $d_{\mathcal{A}}^{N-1}$, where $d_{\mathcal{A}}$ is dimension of a single ancillary system in the collisional strategy.
We can now create a tensor network representing the collisional strategy. In the following we will use functions from qmtensor/creators folder, see Fig. 2. These functions greatly simplify the process by creating the whole structures (for example the density matrix of an MPS) and they make the code more readable. Let us start by defining the necessary spaces:
@@
Tensors of parameter encoding channels can be created using tensor method from the ParamChannel class:
1# Parameter encoding channel:
2ch = par_dephasing(0.75)
@@
We want the control operations to be CPTP maps hence we will use cptp_var creator:
@@
Similarly, as in the case of the input state, the measurement can be created in one of two ways:
@@
To optimize our strategy we need to simply put it into iss_opt function:
1qfi, qfis, tn_opt, status = iss_opt(tn)
where qfi is the optimized QFI, qfis is pre-QFI value per iteration, status tells whether the optimization converged successfully and tn_opt is a copy of tn where each VarTensor is replaced with a ConstTensor storing the result of strategy optimization. To access those values we need to use the names of the required tensor and its spaces:
@@
3# Accessing Choi matrix of the i-th
4# channel:
@@
17# Accessing MPS density matrix
18# i-th element:
@@
Figs. 11 (B) and (C) show the optimized values of the QFI for collisional strategy together with the parallel and adaptive ones for comparison. It turns out that in case of the dephasing noise (B), additional control operations do not help and the results are the same as for the parallel strategy. On the other hand, for the amplitude damping noise (C) the collisional strategy is significantly better than the parallel one and appears to be able to beat the adaptive strategy for $N>100$, under given limited ancillary dimension sizes. This proves that in this case, the optimal strategy found is qualitatively different from both the parallel and the adaptive ones.
@@ 5 Summary
We hope that the presented package will serve as an indispensable tool for anyone, who wants to identify optimal protocols in noise metrology scenarios as well as for those who want to compare the performance of strategies of particular structure with fundamental bounds.
@@
While the package is at the moment focused solely on the single parameter QFI optimization paradigm, it may be in principle developed further to cover Bayesian as well as multi-parameter frameworks. In case of Bayesian framework, though, we do not expect tensor-network methods to work as efficiently as in the QFI case. This is mainly due to the need to use larger ancilla spaces to effectively feed-forward information which is systematically increasing with the growing protocol length, see [22] for more discussion. This would limit applicability of these methods to small scale problems only, and go against the main purpose of this package, which is to provide numerical methods that work efficiently for large-scale quantum metrology problems. For people interested in small-scale Bayesian estimation problems, we can recommend the QuanEstimation package [48], where implementation of some of quantum Bayesian estimation methods can be found.
@@
In general, QuanEstimation package may be regarded as largely complementary to the present one. It provides a wider set of quantum estimation tools, combining both QFI-based and Bayesian methods. The package also has numerous tools dedicated to analyzing and optimizing experimentally relevant quantum control frameworks, as well as some multi-parameter quantum estimation tools. QuanEstimation has optimization procedures that allow to optimize input states in a basic single quantum channel estimation problem, as well as optimize adaptive strategies, with some predefined control structures. This said, efficiency of the proposed algorithms is limited to small scale problems, and even in this case, the optimization procedures employed do not guarantee that the result obtained is indeed optimal.
@@
In case of multi-parameter estimation strategies, one may generalize the ISS and tensor-network methods in a rather straightforward way to optimize the trace of the QFI matrix, as in [49]. However, despite some novel works on this topic [50], a rigorous and at the same time numerically effective optimization of multi-parameter cost that could be applied to study the performance of noisy metrological protocols in the large $N$ limit seems to be out of reach at this moment.
@@
Finally, we should also mention, that we have independently studied and developed similar methods to study the channel discrimination problem [51], results of which will be reported elsewhere.
@@
Acknowledgments. We thank Krzysztof Chabuda for fruitful discussions regarding the inner workings of the TNQMetro package. This work has been supported by the “Quantum Optical Technologies” (FENG.02.01-IP.05-0017/23) project, carried out within the Measure 2.1 International Research Agendas programme of the Foundation for Polish Science co-financed by the European Union under the European Funds for Smart Economy 2021-2027 (FENG), as well as National Science Center (Poland) grant No.2020/37/B/ST2/02134.

### v2

@@ 47 Tsr = x * ket_bra(
51# If not provided, the derivative is
52# zero by default:
53T = ParamChannel(
@@
55)
56
57Lambda_tilde = Phi.link_env(T)
58# or equivalently:
59Lambda_tilde = Phi * T
@@
4from qmetro import iss_tnet_adaptive_qfi
5
6N = 3
7d_a = 2
8# QFI for the adaptive strategy and
9# correlated noise:
10qfi, *_ = iss_tnet_adaptive_qfi(
11 Lambda_tilde, N, d_a
12)
13
14# Input state of the environment is by
15# default a maximally mixed state but
16# can be specified using env_inp_state
17# parameter:
18rhoE = identity(2) / 2
19qfi, *_ = iss_tnet_adaptive_qfi(
20 Lambda_tilde, N, d_a,
21 env_inp_state=rhoE
22)
Bounds in this case can be computed using ad_bounds_correlated and ad_asym_bound_correlated. In contrast to the bounds for the uncorrelated case these bounds are not necessarily tight asymptotically. They are computed by upper-bounding QFI for adaptive strategy where every $m$-th control operation has access to the environment, for some integer $m$. This leads to a leakage of information from the environment, which weakens the tightness of the bound. Naturally, larger value of $m$ makes the bound tighter but at the same time it exponentially increases the execution time. Realistically, one can compute the bound for $m\lesssim 5$ in case of qubit channels. An exemplary code to compute the bound is given below:
@@ 20 Lambda_tilde, N, d_a,
3from qmetro import ad_bounds_correlated
4
5N = 10
6m = 3
7ns, bound = ad_bounds_correlated(
8 Lambda_tilde, N, m
9)
[…]
Note that the construction (51), (52) is not unique and correlations (50) can be modeled using different $\tilde{\Lambda}_{\theta}$. For example a quantum channel:
$|,\mathrm{t}:\mathcal{L}(\mathcal{E})\ni\rho\mapsto\sum_{s,r}t_{sr}\rho\,t_{sr}^{\dagger}\in\mathcal{L}(\mathcal{E}),\\ t_{sr}=\sqrt{\frac{1+sr\sqrt{c}}{2}}\ket{s}\bra{r}$, (53)
@@ 8 Lambda_tilde, N, m
Values of the QFI for anti-correlated (c=0.75c=-0.75) dephasing case are shown in Figs.7 (B) and (C). Note, that the presence of (anti)correlations significantly increases the obtained QFI which becomes even larger then the upper-bounds for uncorrelated case. Interestingly, we can see that the adaptive strategy with $d_{\mathcal{A}}=2$ is optimal for $N\leq 5$, unlike in the uncorrelated dephasing case, see Fig. 6(B).
@@ 4 Advanced package usage—optimization of strategies with arbitrary structures
Figure 9: Diagram of a strategy with arbitrary structure that can be optimized using iss_opt. The input state, $\rho_{0}$, might consist of MPSs and/or states on a full Hilbert space. The input state goes through an arbitrary arrangement of parametrized channels, control operations and combs. The requirement is that each operation’s output can be connected only to the input of a different operation such that there are no causal cycles. Finally, the state obtained after application of each operation, $\rho_{\theta}$, is measured. The measurement is specified by the form of the pre-SLD matrix $\mathfrak{L}$ which can be an MPO and/or a matrix on a full Hilbert space. States, control operations, and combs can be marked as variable or constant, the pre-SLD must be variable and the parametrized channels must be constant.
In addition to the functions presented in Sec. 3, the QMetro++ allows the user to define their own strategy and then optimize it using the ISS algorithm. Strategies are defined via a straightforward-to-use symbolic programming framework. The user creates a tensor network representing their strategy where some of the nodes are marked as constants and others are marked as variables to be optimized over—let us call the set of variables $\mathcal{V}$. This strategy is then plugged into the function iss_opt which, based on the provided data, constructs the pre-QFI:
@@
In Sec. 4.1 we explain the classes of QMetro++ tensors (placed in qmtensor folder, see Fig. 2) which are the building blocks of strategies. Then in Sec. 4.2 we show how such a strategy can be created using the example of the collisional metrological model.
@@ 4.1 QMetro++ tensors
  • parameter-encoding channels,
@@
connected by their physical and bond spaces, as in the diagram in Fig. 9. Some input states, control operations, and quantum combs as well as all measurements, are marked as variables of the pre-QFI function (54) and they will be optimized by the ISS algorithm. Each non-measurement node is represented using its tensor representation (see a detailed explanation in the Appendix B) and measurements are represented via a tensor representation of the pre-SLD matrix $\mathfrak{L}$.
Figure 10: UML diagram of classes used for creation of strategies with their most important attributes and methods. Subclasses inherit all attributes and methods of their parent class. Instances of GeneralTensor and its subclasses can access a SpaceDict instance through their sdict attribute. Two tensors can be contracted with each other only when they have the same sdict attribute.
In practice, this means that each node of the network has to be initialized as an object of a QMetro++ tensor class (see Fig. 10). There are five kinds of tensor classes:
@@
  • ParamTensor - a subclass of ConstTensor which has an additional attribute defining its derivative over the measured parameter $\theta$. It is meant to represent parameter-encoding channels $\Lambda_{\theta}$.
  • VarTensor - a class representing tensors that are optimized by the ISS algorithm. Members of this class store information specifying how they should be initialized and optimized (for example that they are CPTP channels, quantum combs or pre-SLD matrices) but they do not have any value—they are only symbols.
@@
10# Calling sd[’a’] will return
11# the dimension of space ’a’.
@@
15sd.set_bond(’x’, 4)
16print(sd[’x’]) # Prints 4.
@@
20print(sd.irange[’x’]) # Prints 4.
Note that the range of the physical index is equal to the square of the dimension of the corresponding space. This is because physical indices are created by joining two indices of a CJ operator. However, this is not the case for bond indices which are not related to any Hilbert spaces. See detailed discussion in Appendix B).
@@
Variable tensors require only spaces. Their type is derived from additional attributes:
@@
14# > element of a density matrix of
15# an MPS:
16vt1 = VarTensor(
@@
19)
20# > CPTP map of a channel ’a’->’b’:
21vt1 = VarTensor(
@@
3channel = par_dephasing(0.75)
4choi = channel.choi()
@@
While contraction of two ConstTensors simply creates a new ConstTensor this is not the case when one of the contracted tensors is a VarTensor because it does not have any value. Contraction with a VarTensor creates a TensorNetwork consisting of the two tensors:
@@
Similarly, a contraction with a TensorNetwork creates another TensorNetwork. The results of all possible contraction combinations are presented in Table 2.
@@
Table 2: Resulting type of a contraction of various tensor pairs: GT–GeneralTensor, CT–ConstTensor, PT–ParamTensor, VT– VarTensor, TN–TensorNetwork. Contraction with generalized tensor raises an error because instances of this class are not meant to be used in calculations. When ParamTensor is contracted with a ConstTensor or a ParamTensor the derivative of the result is computed using the Leibniz (chain) rule.
@@ 4.2 Collisional strategy
Figure 11: (A) Diagram of a collisional strategy with multiple parametrized channels $\Lambda_{\theta}$ and ancillas $\mathcal{A}_{i}$. Values of QFI normalized by the number of channel uses, $N$, for: (B) dephasing ($p=0.75$), (C) amplitude damping ($p=0.75$) and different strategies: adaptive ($d_{\mathcal{A}}=2$) - blue dashed line, parallel ($d_{\mathcal{A}}=2,\;r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}=2$) - red dotted line, collisional ($d_{\mathcal{A}}=2,\;r_{\mathrm{MPS}}=\sqrt{r_{\mathfrak{L}}}=2$) - green dash-dotted line. Fluctuations for larger NN are due to numerical instabilities.
We will describe the creation of a custom strategy using the example of a collisional strategy depicted in Fig. 11(A). Physically, we can imagine this scheme to represent a situation where multiple particles, potentially entangled with each other initially, approach the sensing system one by one. They interact locally with the system via control operations $\mathrm{C}_{i}$, carrying away information about the parameter while at the same time causing some back-action on the sensing system itself.
This strategy may be viewed as a hybrid of the parallel and the adaptive strategies where each control operation $\mathrm{C}_{i}$ has its own ancilla which is measured immediately after the action of $\mathrm{C}_{i}$. Notice that if all $\mathrm{C}_{i}$ were set to SWAP gates one would regain the parallel strategy. On the other hand, this strategy does not allow control operations to communicate with each other like in the most general adaptive strategy. Therefore the QFI for this strategy must be at least as good as for the parallel strategy but in general will be worse than for the adaptive strategy with an ancilla dimension $d_{\mathcal{A}}^{N-1}$, where $d_{\mathcal{A}}$ is the dimension of a single ancillary system in the collisional strategy.
We can now create a tensor network representing the collisional strategy. In the following we will use functions from the qmtensor/creators folder, see Fig. 2. These functions greatly simplify the process by creating entire structures (for example the density matrix of an MPS) and they make the code more readable. Let us start by defining the necessary spaces:
@@
Tensors of parameter-encoding channels can be created using tensor method from the ParamChannel class:
1# Parameter-encoding channel:
2channel = par_dephasing(0.75)
@@
We want the control operations to be CPTP maps, hence we will use cptp_var creator:
@@
Similarly to the input state, the measurement can be created in one of two ways:
@@
To optimize our strategy, we need to simply put it into iss_opt function:
1qfi, qfis, tnet_opt, status = iss_opt(tnet)
where qfi is the optimized QFI, qfis is a list of pre-QFI values per iteration, status tells whether the optimization converged successfully and tnet_opt is a copy of tnet where each VarTensor is replaced with a ConstTensor storing the result of the strategy optimization. To access those values we need to use the names of the required tensor and its spaces:
@@
3# Accessing the CJ matrix of the i-th
4# control operation:
@@
17# Accessing the i-th element
18# of the MPS density matrix:
@@
Figs. 11 (B) and (C) show the optimized values of the QFI for collisional strategy together with the parallel and the adaptive ones for comparison. It turns out that in the case of the dephasing noise (B), additional control operations do not help and the results are the same as for the parallel strategy. On the other hand, for the amplitude damping noise (C) the collisional strategy is significantly better than the parallel one and appears to be able to beat the adaptive strategy for $N>100$, under the given limited ancillary dimensions. This proves that in this case, the optimal strategy found is qualitatively different from both the parallel and the adaptive ones.
@@ 5 Summary
We hope that the presented package will serve as an indispensable tool for anyone who wants to identify optimal protocols in noisy metrology scenarios as well as for those who want to compare the performance of strategies of a particular structure with fundamental bounds.
@@
While the package is at the moment focused solely on the single-parameter QFI optimization paradigm, it may be in principle developed further to cover Bayesian as well as multi-parameter frameworks. In case of Bayesian framework, though, we do not expect tensor-network methods to work as efficiently as in the QFI case. This is mainly due to the need to use larger ancilla spaces to effectively feed-forward information that systematically increases with the protocol length, see [31] for further discussion. This would limit applicability of these methods to small-scale problems only, and go against the main purpose of this package, which is to provide numerical methods that work efficiently for large-scale quantum metrology problems. For people interested in small-scale Bayesian estimation problems, we can recommend the QuanEstimation package [49], where implementations of some of the quantum Bayesian estimation methods can be found.
@@
In general, the QuanEstimation package may be regarded as largely complementary to the present one. It provides a wider set of quantum estimation tools, combining both QFI-based and Bayesian methods. The package also has numerous tools dedicated to analyzing and optimizing experimentally relevant quantum control frameworks, as well as some multi-parameter quantum estimation tools. QuanEstimation has optimization procedures that allow to optimize input states in a basic single quantum channel estimation problem, as well as optimize adaptive strategies, with some predefined control structures. That said, efficiency of the proposed algorithms is limited to small-scale problems, and even in this case, the optimization procedures employed do not guarantee that the result obtained is indeed optimal.
@@
In case of multi-parameter estimation strategies, one may generalize the ISS and tensor-network methods in a rather straightforward way to optimize the trace of the QFI matrix, as in [1]. However, despite some novel works on this topic [25], a rigorous and at the same time numerically effective optimization of a multi-parameter cost that could be applied to study the performance of noisy metrological protocols in the large $N$ limit seems to be out of reach at this stage.
@@
Finally, we should also mention that we have independently studied and developed similar methods to study the channel discrimination problem [4], results of which will be reported elsewhere.
@@
Acknowledgments. We thank Krzysztof Chabuda for fruitful discussions regarding the inner workings of the TNQMetro package. This work has been supported by the “Quantum Optical Technologies” (FENG.02.01-IP.05-0017/23) project, carried out within the Measure 2.1 International Research Agendas programme of the Foundation for Polish Science co-financed by the European Union under the European Funds for Smart Economy 2021-2027 (FENG), as well as National Science Center (Poland) grant No.2020/37/B/ST2/02134. SK was supported by the Foundation for Polish Science (FNP).